Suppose that you invest $1800 in an account that earns simple interest at an APR of 5.2%. Determine the accumulated balance after 5 years.
Answer by stanbon(75887)

Suppose that you invest $1800 in an account that earns simple interest at an APR of 5.2%. Determine the accumulated balance after 5 years.
---------------
A(t) = P(1+r*t)
--------------------
A(5) = 1800(1+0.052*5) = $2268
-------------
Cheers,
Stan H.
-----------